OpenStudy (anonymous):

Solve for x. 2+1/x-3/x^2=0

OpenStudy (anonymous):

divide thru by the lcd to get rid of the fraction lcd = x^2 this leaves a quadratic equation to be solved

OpenStudy (anonymous):

this is 2x^2 + x -3 = 0 which can be factorized

OpenStudy (anonymous):

Divide through? Or multiply through?

OpenStudy (anonymous):

2+(1)/(x)-(3)/(x^(2))=0 Find the LCD (least common denominator) of 2+(1)/(x)-(3)/(x^(2))+0. Least common denominator: x^(2) Multiply each term in the equation by x^(2) in order to remove all the denominators from the equation. 2*x^(2)+(1)/(x)*x^(2)-(3)/(x^(2))*x^(2)=0*x^(2) Cancel the common factor of x in the denominator of the first term (1)/(x) and the second term x^(2). 2*x^(2)+(1)/(<X>x<x>)*x^(<X>2<x>)-(3)/(x^(2))*x^(2)=0*x^(2) Reduce the expression by removing the common factor of x in the denominator of the first term (1)/(x) and the second term x^(2). 2*x^(2)+1*x-(3)/(x^(2))*x^(2)=0*x^(2) Cancel the common factor of x^(2) in the denominator of the first term -(3)/(x^(2)) and the second term x^(2). 2*x^(2)+1*x-(3)/(x^(<X>2<x>^(0)))*x^(<X>2<x>^(0))=0*x^(2) Reduce the expression by removing the common factor of x^(2) in the denominator of the first term -(3)/(x^(2)) and the second term x^(2). 2*x^(2)+1*x-3*1=0*x^(2) Multiply 2 by x^(2) to get 2x^(2). 2x^(2)+1*x-3*1=0*x^(2) Multiply 1 by x to get x. 2x^(2)+x-3*1=0*x^(2) Multiply -3 by 1 to get -3. 2x^(2)+x-3=0*x^(2) Multiply 0 by x^(2) to get 0. 2x^(2)+x-3=0 For a polynomial of the form ax^(2)+bx+c, find two factors of a*c (-6) that add up to b (1).In this problem (3)/(2)*-1=-(3)/(2) (which is (c)/(a)) and (3)/(2)-1=(1)/(2) (which is ((b)/(a)) , so insert (3)/(2) as the right hand term of one factor and -1 as the right-hand term of the other factor. (x+(3)/(2))(x-1)=0 Remove the fraction by multiplying the first term of the factor by the denominator of the second term. (2x+3)(x-1)=0 Set each of the factors of the left-hand side of the equation equal to 0. 2x+3=0_x-1=0 Since 3 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 3 from both sides. 2x=-3_x-1=0 Divide each term in the equation by 2. (2x)/(2)=-(3)/(2)_x-1=0 Cancel the common factor of 2 in (2x)/(2). (<X>2<x>x)/(<X>2<x>)=-(3)/(2)_x-1=0 Remove the common factors that were cancelled out. x=-(3)/(2)_x-1=0 Set each of the factors of the left-hand side of the equation equal to 0. x=-(3)/(2)_x-1=0 Since -1 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 1 to both sides. x=-(3)/(2)_x=1 The complete solution is the set of the individual solutions. x=-(3)/(2),1

OpenStudy (radar):

\[2 + 1/x-3/x ^{2}=0\] Multiply all terms by x^2 \[2x ^{2}+x-3=0\] Factor (x-1)(2x+3)=0 solve by setting each factor to 0 x-1=0, x=1 2x+3=0 2x=-3 x=-3/2 This is the same solution as above (but abbreviated)

OpenStudy (anonymous):

Now I AM GOING TO GIVE A SIMPLE SULOTION (the above is too long). Here: Make 2 + 1/x -3/x^2 = 0 to (by multiplying x^2) 2x^2 + x -3 =0/(1/x^2)=0 Then factorize it to:(x-1)(2x+3)=0 so x-1=0 and 2x+3=0 Then x = 1 or x=-3/2